OpenStudy (anonymous):

solve: 4(x - 2) - 10x = -40

OpenStudy (anonymous):

Use the distributive property first. \(a(b + c) = ab + ac\) So in this case:\[4(x - 2) = 10x = -40\]\[(4x - 8) - 10x = -40\]

OpenStudy (austinl):

Combine like terms, get constants on one side, and variables on the other.